# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 3.1 Fibonacci Preliminaries
The resulting “carryless” coding discipline and band separation refine this tradition. The additive Fibonacci recurrence yields a canonical representation with explicit inversion, while avoiding interference between adjacent components. This separation permits controlled aggregation of constraints, and the Diophantine degree cost remains quadratic throughout the encoding layer. See Rosko [2026] for a self-contained treatment of the construction and its formal properties.
Definition 3.1 (Fibonacci canonicalization).
@@
$F_{0}=0,\quad F_{1}=1,\quad F_{n+2}=F_{n+1}+F_{n}\quad\text{for all }n\geq 0.$ (3.40)
Lemma 3.2 (Fibonacci monotonicity).
@@
$F_{n+1}\geq F_{n},\quad F_{n+2}>F_{n}.$ (3.41)
Induction coincides with the recursive definition. ∎
Theorem 3.3 (Zeckendorf’s Theorem).
@@
$n=\sum_{k\in S}F_{k},\quad S\subseteq\{2,3,4,\ldots\}$ (3.42)
@@
By Zeckendorf [1972]. The non-adjacency constraint is preserved by the monotonic gap property $F_{k+2}>F_{k}$. Assume two distinct representations for $n$. Let $k_{\max}$ be the largest index where they differ. Without loss of generality, suppose the first representation includes $F_{k_{\max}}$ and the second does not. The Fibonacci identity
$\sum_{i=2}^{k_{\max}-2}F_{i}=F_{k_{\max}}-1$ (3.43)
@@ 3.2 Carryless Pairing
Intuition.
The Zeckendorf Representation assigns to each $n$ a canonical support set $Z(n)\subseteq\mathbb{N}$. We construct an injective pairing by embedding the supports of $x$ and $y$ into disjoint frequency bands within a combined Zeckendorf expansion. The even-indexed Fibonacci numbers encode $x$; the odd-indexed numbers, shifted beyond the range of $x$’s encoding, encode $y$. No arithmetic carries occur and composition is achieved by set-theoretic union of supports. For a presentation with consult either Rosko [2026] or the repository:
@@
Theorem 3.4 (Carryless Pairing).
Define $Z(0):=\varnothing$. For $x,y\in\mathbb{N}$, set
$E(x)$ :={ $:=\{\,2k:k\in Z(x)\,\},$ (3.44)
$r(x)$ $:=\min\{\,m\geq 2:F_{m}>x\,\},$ (3.45)
$B(x)$ $:=2r(x),$ (3.46)
$O(x,y)$ $:=\{\,B(x)+(2j-1):j\in Z(y)\,\}.$ (3.47)
Define the pairing
$\mathrm{pair}(x,y)\;:=\;F_{2}\;+\;\sum_{i\in E(x)\cup O(x,y)}F_{i}.$ (3.48)
@@
$\max E(x)<B(x)\leq\min O(x,y)-3,$ (3.49)
and $\mathrm{pair}$ is injective. Moreover there are primitive recursive projections $\mathrm{hd},\mathrm{tl}$ such that
$u>0\;\to\;\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u,\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0,$ (3.50)
and $\mathrm{unpair}(u)=(\mathrm{hd}(u),\mathrm{tl}(u))$ is the inverse of $\mathrm{pair}$ on $\mathbb{N}^{2}$.
By induction over uniquness, construction of $B(x)=2r(x)$, and rank definition. Band separation ensures the union $E(x)\cup O(x,y)$ is a valid Zeckendorf support. By uniqueness (Theorem 3.3), the support is recoverable; the even and odd bands are then separated and decoded independently. ∎
Definition 3.5 (List encoding).
@@
$\mathrm{code}_{\mathrm{list}}([])$ $:=0,$ (3.51)
$\mathrm{code}_{\mathrm{list}}(x_{1},\ldots,x_{n})$ $:=\mathrm{pair}(x_{1},\mathrm{code}_{\mathrm{list}}(x_{2},\ldots,x_{n})).$ (3.52)
The function $\mathrm{code}_{\mathrm{list}}$ inherits injectivity from Theorem 3.4.
Remark.
@@ 3.3 Carryless Sequence Indexing
Intuition.
Finite proof traces are encoded directly by nested Carryless Pairing. The same pairing infrastructure used for formula coding also provides line access: the $i$-th entry is obtained by iterating the tail projection $i$ times and reading the head. This removes any dependence on modular sequence decoding while keeping extraction primitive recursive and degree-local.
Definition 3.6 (Head and tail projections).
Let $\mathrm{pair}:\mathbb{N}^{2}\to\mathbb{N}$ be the carryless pairing of Theorem 3.4. Define primitive recursive projections
$\mathrm{hd},\mathrm{tl}:\mathbb{N}\to\mathbb{N}$ (3.53)
@@
$\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u\;\;(u>0),\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0.$ (3.54)
Definition 3.7 (Indexed extractor).
@@
$\mathrm{tail}^{0}(s)$ $:=s,$ (3.55)
$\mathrm{tail}^{i+1}(s)$ $:=\mathrm{tl}(\mathrm{tail}^{i}(s)),$ (3.56)
$\mathrm{line}(s,i)$ $:=\mathrm{hd}(\mathrm{tail}^{i}(s)).$ (3.57)
Theorem 3.8 (Carryless sequence representation).
@@
$\mathrm{line}(s,i)=a_{i}\quad\text{for all }i<n.$ (3.58)
Take $s:=\mathrm{code}_{\mathrm{list}}(a_{0},\ldots,a_{n-1})$ from Definition 3.5. The claim follows by induction on $i$, using invertibility of $\mathrm{pair}$ (Theorem 3.4). ∎
Definition 3.9 (Bounded extraction witnesses).
@@
$t_{i,0},\ldots,t_{i,i+1},\quad h_{i,0},\ldots,h_{i,i}.$ (3.59)
@@
$t_{i,0}$ $=s,$ (3.60)
$t_{i,r}$ $=\mathrm{pair}(h_{i,r},t_{i,r+1})\quad(0\leq r\leq i).$ (3.61)
@@ 3.4 Coding of Syntax
Definition 3.10 (Syntactic and proof predicates).
@@
$\operatorname{Sent}(u)$ $\;:\equiv\;\operatorname{Form}(u)\;\wedge\;u\text{ codes a sentence},$ (3.63)
$\operatorname{Proof}_{\mathsf{RA}}(p,u)$ $\;:\equiv\;p\text{ codes a valid $\mathsf{RA}$-proof coded by }u,$ (3.64)
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;:\equiv\;\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u).$ (3.65)
Observe how $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ holds when $p$ codes a finite sequence such that $\varphi_{n-1}$ is the sentence coded by $u$, and for each $i<n$, either $\varphi_{i}$ is an instance of an axiom schema of $\mathsf{RA}$, or there exist $j,k<i$ with $\varphi_{k}\equiv\varphi_{j}\to\varphi_{i}$ by Modus ponens.
Lemma 3.11 (Primitive recursiveness).
The predicates $\operatorname{Form}$, $\operatorname{Sent}$, and $\operatorname{Proof}_{\mathsf{RA}}$ are primitive recursive, consequently, $\operatorname{Prov}_{\mathsf{RA}}$ is $\Sigma_{1}$.
By definition. Each predicate is decided by bounded inspection of the input code. $\operatorname{Form}$ and $\operatorname{Sent}$ are checked by structural induction on code depth. $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ is checked by scanning the sequence coded by $p$: at each line, either an axiom-schema match is verified by pattern matching, or a Modus ponens witness is found by search bounded by the line index. All checks are bounded computations, hence primitive recursive. The $\Sigma_{1}$-classification of $\operatorname{Prov}_{\mathsf{RA}}$ follows from the existential quantification over $p$. ∎
Definition 3.12 (Bridge translation).
Let $\operatorname{tr}:\mathcal{L}_{\mathsf{RA}}\to\mathcal{L}_{\bot,\to}$ be the primitive recursive encoding used by the kernel checker $\mathrm{check}_{\to}(p,u)$ on the implicational language. The bridge is captured by two numeralwise implications:
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;\Longrightarrow\;\mathsf{RA}\vdash\exists p\,\mathrm{check}_{\to}(p,\operatorname{tr}(\bar{u})),$ (3.66)
@@
That is, provable sentences have kernel proofs (formalizable in $\mathsf{RA}$), and any kernel proof witnesses truth of the corresponding $\mathsf{RA}$ sentence in the standard model. No global biconditional inside $\mathsf{RA}$ is assumed.
@@ 3.5 Fixed-Point Construction
Theorem 3.13 (Derivability Conditions).
@@
$\textsf{D1}:\;$ $\text{If}\;\mathsf{RA}\vdash\varphi,\;\text{then}\;\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner).$ (3.68)
$\textsf{D2}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\urcorner\bigr).$ (3.69)
$\textsf{D3}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\to\psi\urcorner)\to\bigl(\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\psi\urcorner)\bigr).$ (3.70)
These conditions are required for Gödel’s 2nd as applied in Corollary 5.7.
$\mathsf{D1}$ is immediate: a proof of $\varphi$ in $\mathsf{RA}$ is a concrete finite sequence whose code witnesses $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)$. (D2) and (D3) are standard consequences of the primitive recursiveness of $\operatorname{Proof}_{\mathsf{RA}}$ (Lemma 3.11) and the representability of primitive recursive functions in $\mathsf{RA}$ (Proposition 1.10): $\mathsf{RA}$ can internally simulate concatenation of proof sequences and verify that the result is again a valid proof. See Boolos et al. [2012]. ∎
Theorem 3.14 (Diagonal Lemma).
@@
$\mathsf{RA}\;\vdash\;u\;\leftrightarrow\;\varphi(\ulcorner u\urcorner).$ (3.71)
Define the substitution function $\mathrm{sub}(e,n)$ to be the code of the sentence obtained by substituting the numeral $\overline{n}=S^{n}(0)$ for the free variable in the formula coded by $e$. The function $\mathrm{sub}$ is primitive recursive, it acts by bounded structural replacement in the formula tree, and is therefore representable in $\mathsf{RA}$ by Proposition 1.10. Let
$\psi(x):\equiv\varphi(\mathrm{sub}(x,x))\quad\text{with}\quad e:=\ulcorner\psi\urcorner$ (3.72)
@@
$u\;\equiv\;\psi(\overline{e})\;\equiv\;\varphi\bigl(\mathrm{sub}(\overline{e},\overline{e})\bigr)\;\equiv\;\varphi(\ulcorner u\urcorner),$ (3.73)
@@
4 From Proof Checking to Polynomial Constraints
@@ 4.1 The Hilbert–Ackermann Kernel
Intuition.
We fix a minimal proof system for propositional logic in the style of Hilbert and Ackermann [1938], with implication and falsity as primitive connectives. A proof is a finite sequence of formulas, each either an axiom instance or a consequence of Modus ponens from earlier lines, concluding at the target formula. A total boolean checker decides correctness of any proposed proof; soundness (Theorem 4.5) confirms it is faithful to derivability. Being primitive recursive, the checker arithmetizes directly via the encoding mechanisms of the preceding sections. We then reflect each of its verification conditions into polynomial equalities over $\mathbb{N}$, obtaining a finite constraint system whose solvability is equivalent to checker acceptance (Theorem 4.19) with every emitted equation of degree at most $3$ (Proposition 4.20).
Definition 4.1 (Formula syntax).
@@
$\mathsf{Form}\;::=\;\bot\mid(\varphi\to\psi),$ (4.74)
where $\bot$ denotes falsity and $\varphi\to\psi$ denotes implication. The mechanized kernel ($\mathtt{R04}$) uses exactly these two constructors; propositional variables are not needed, since the axiom schemas already quantify over all formulas.
Definition 4.2 (Axiom schemas).
The proof system admits three axiom schemas [Troelstra and Schwichtenberg, 2000]. The schemas K and S are the logical counterparts of the combinators $\mathbf{K}$ and $\mathbf{S}$. For all formulas $\varphi,\psi,\chi\in\mathsf{Form}$:
@@
    $\big(\varphi\to(\psi\to\chi))\to((\varphi\to\psi)\to(\varphi\to\chi)\big).$ (4.76)
@@
    $\bot\to\varphi.$ (4.77)
Definition 4.3 (Proof sequence).
@@
$p=(\varphi_{0},\ldots,\varphi_{n-1})$ (4.78)
@@
      Axiom: $\varphi_{i}$ is an instance of one of the axiom schemas (Definition 4.2).
@@
      $\varphi_{k}=(\varphi_{j}\to\varphi_{i}).$ (4.79)
@@
    $\varphi_{n-1}=\varphi.$ (4.80)
Definition 4.4 (Proof checker).
@@
$\mathrm{check}:\mathsf{ProofSeq}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\}$ (4.81)
that returns $\mathrm{true}$ if and only if $p$ satisfies the conditions of Definition 4.3 with target $\varphi$. The search for Modus ponens witnesses at each line $i$ is bounded by $i$, so $\mathrm{check}$ is total and primitive recursive.
Theorem 4.5 (Kernel soundness).

### v7

@@ 3.1 Fibonacci Preliminaries
The resulting “carryless” discipline and band separation refine this tradition. The additive Fibonacci recurrence yields a canonical representation with explicit inversion, while avoiding interference between adjacent components. This separation permits controlled aggregation of constraints, and the Diophantine degree cost remains quadratic throughout the encoding layer.
Definition 3.1.2 (Fibonacci canonicalization).
@@
$F_{0}=0,\quad F_{1}=1,\quad F_{n+2}=F_{n+1}+F_{n}\quad\text{for all }n\geq 0.$ (54)
Lemma 3.1.3 (Fibonacci monotonicity).
@@
$F_{n+1}\geq F_{n},\quad F_{n+2}>F_{n}.$ (55)
Immediate. Induction coincides with the recursive definition. ∎
Theorem 3.1.4 (Zeckendorf’s Theorem).
@@
$n=\sum_{k\in S}F_{k},\quad S\subseteq\{2,3,4,\ldots\}$ (56)
@@
By induction [Zeckendorf, 1972]. The non-adjacency constraint is preserved by the monotonic gap property $F_{k+2}>F_{k}$. Assume two distinct representations for $n$. Let $k_{\max}$ be the largest index where they differ. Without loss of generality, suppose the first representation includes $F_{k_{\max}}$ and the second does not. The Fibonacci identity
$\sum_{i=2}^{k_{\max}-2}F_{i}=F_{k_{\max}}-2$ (57)
@@ 3.2 Carryless Pairing
Remark 3.2.1.
The Zeckendorf Representation assigns to each $n$ a canonical support set $Z(n)\subseteq\mathbb{N}$. We construct an injective pairing by embedding the supports of $x$ and $y$ into disjoint frequency bands within a combined Zeckendorf expansion. The even-indexed Fibonacci numbers encode $x$; the odd-indexed numbers, shifted beyond the range of $x$’s encoding, encode $y$. No arithmetic carries occur and composition is achieved by set-theoretic union of supports. For a presentation, consult either Rosko [2026] or the repository:
@@
Theorem 3.2.2 (Carryless Pairing).
Over $Z(0):=\varnothing$ and $x,y\in\mathbb{N}$, set
$E(x)$ :={ $:=\{\,2k:k\in Z(x)\,\},$ (58)
$r(x)$ $:=\min\{\,m\geq 2:F_{m}>x\,\},$ (59)
$B(x)$ $:=2r(x),$ (60)
$O(x,y)$ $:=\{\,B(x)+(2j-1):j\in Z(y)\,\}.$ (61)
Define a pairing function,
$\mathrm{pair}(x,y)\;:=\;F_{2}\;+\;\!\!\!\sum_{i\in E(x)\cup O(x,y)}\!\!\!F_{i}.$ (62)
@@
$\max E(x)<B(x)\leq\min O(x,y)-3,$ (63)
and $\mathrm{pair}(\cdot)$ is injective. Moreover there are primitive recursive projections $\mathrm{hd},\mathrm{tl}$ such that
$u>0\;\to\;\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u,\quad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0,$ (64)
and
$\mathrm{unpair}(u)=(\mathrm{hd}(u),\mathrm{tl}(u))$ (65)
is the inverse of $\mathrm{pair}$ on $\mathbb{N}^{2}$.
By construction. Band separation, $B(x)=2r(x),$ ensures the union
$E(x)\cup O(x,y)$ (66)
is a valid Zeckendorf support. By uniqueness (Theorem 3.1.4), the support is recoverable; the even and odd bands are then separated and decoded independently. ∎
Definition 3.2.3 (List encoding).
@@
$\mathrm{code}_{\mathrm{list}}([])$ $:=0,$ (67)
$\mathrm{code}_{\mathrm{list}}(x_{1},\ldots,x_{n})$ $:=\mathrm{pair}(x_{1},\mathrm{code}_{\mathrm{list}}(x_{2},\ldots,x_{n})).$ (68)
The function $\mathrm{code}_{\mathrm{list}}$ inherits injectivity from Theorem 3.2.2.
Remark 3.2.4.
@@ 3.3 Carryless Sequence Indexing
Remark 3.3.1.
Finite proof traces are encoded directly by Carryless Pairing. The same infrastructure used for formula coding also provides line access: the $i$-th entry is obtained by iterating the tail projection $i$ times and reading the head. This removes any dependence on modular sequence decoding while keeping extraction primitive recursive and degree-local.
Definition 3.3.2 (Head and tail projections).
Let $\mathrm{pair}:\mathbb{N}^{2}\to\mathbb{N}$ be the carryless pairing of Theorem 3.2.2. Define primitive recursive projections
$\mathrm{hd},\mathrm{tl}:\mathbb{N}\to\mathbb{N}$ (69)
@@
$\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u\;\;(u>0),\quad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0.$ (70)
Definition 3.3.3 (Indexed extractor).
@@
$\mathrm{tail}^{0}(s)$ $:=s,$ (71)
$\mathrm{tail}^{i+1}(s)$ $:=\mathrm{tl}(\mathrm{tail}^{i}(s)),$ (72)
$\mathrm{line}(s,i)$ $:=\mathrm{hd}(\mathrm{tail}^{i}(s)).$ (73)
Theorem 3.3.4 (Carryless sequence representation).
@@
$\mathrm{line}(s,i)=a_{i}\quad\text{for all }i<n.$ (74)
By induction over $s:=\mathrm{code}_{\mathrm{list}}(a_{0},\ldots,a_{n-1})$ from Definition 3.2.3. The claim follows by induction on $i$, using invertibility of $\mathrm{pair}$ (Theorem 3.2.2). ∎
Definition 3.3.5 (Bounded extraction witnesses).
@@
$t_{i,0},\ldots,t_{i,i+1},\quad h_{i,0},\ldots,h_{i,i}.$ (75)
@@
$t_{i,0}$ $=s,$ (76)
$t_{i,r}$ $=\mathrm{pair}(h_{i,r},t_{i,r+1})\quad(0\leq r\leq i).$ (77)
@@ 3.4 Coding of Syntax
Definition 3.4.2 (Syntactic and proof predicates).
@@
$\operatorname{Sent}(u)$ $\;:\equiv\;\operatorname{Form}(u)\;\wedge\;u\text{ codes a sentence},$ (79)
@@
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;:\equiv\;\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u).$ (81)
Observe how $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ holds when $p$ codes a finite sequence such that $\varphi_{n-1}$ is the sentence coded by $u$, and for each $i<n$, either $\varphi_{i}$ is an instance of an axiom schema of $\mathsf{RA}$, or there exist $j,k<i$ with
$\varphi_{k}\equiv\varphi_{j}\to\varphi_{i}$ (82)
by modus ponens.
Lemma 3.4.3 (Primitive recursiveness).
The predicates $\operatorname{Form}$, $\operatorname{Sent}$, and $\operatorname{Proof}_{\mathsf{RA}}$ are primitive recursive; consequently, $\operatorname{Prov}_{\mathsf{RA}}$ is $\Sigma_{1}$.
Immediate. Each predicate is decided by bounded inspection of the input code. $\operatorname{Form}$ and $\operatorname{Sent}$ are checked by structural induction on code depth. $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ is checked by scanning the sequence coded by $p$: at each line, either an axiom-schema match is verified by pattern matching, or a modus ponens witness is found by search bounded by the line index. All checks are bounded computations, hence primitive recursive. The $\Sigma_{1}$-classification of $\operatorname{Prov}_{\mathsf{RA}}$ follows from the existential quantification over $p$. ∎
Definition 3.4.4 (Bridge translation).
Let $\operatorname{tr}:\mathcal{L}_{\mathsf{RA}}\to\mathsf{Form}$ be the primitive recursive encoding of $\mathcal{L}_{\mathsf{RA}}$-sentences as elements of $\mathsf{Form}$ (Definition 4.1.2). Since the kernel checker $\mathrm{check}(p,u)$ validates proofs in the Hilbert system of Definition 4.1.3—which includes both the logical schemas K, S, EFQ and the non-logical axioms of $\mathsf{RA}$—its acceptance set coincides with $\mathsf{RA}$-provability on $\operatorname{tr}$-images. Concretely, the bridge is captured by two numeralwise implications:
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;\Longrightarrow\;\mathsf{RA}\vdash\exists p\,\mathrm{check}(p,\operatorname{tr}(\bar{u})),$ (83)
[…]
No global biconditional inside $\mathsf{RA}$ is assumed.
@@ 3.5 Fixed-Point Construction
Theorem 3.5.1 (Derivability Conditions).
@@
$\textsf{D1}:\;$ $\text{If}\;\mathsf{RA}\vdash\varphi,\;\text{then}\;\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner).$ (85)
$\textsf{D2}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\urcorner\bigr).$ (86)
$\textsf{D3}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\to\psi\urcorner)\to\bigl(\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\psi\urcorner)\bigr).$ (87)
These conditions are required for Gödel’s 2nd as applied in Corollary 5.3.5.
Immediate. $\mathsf{D1}$: a proof of $\varphi$ in $\mathsf{RA}$ is a concrete finite sequence whose code witnesses $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)$. (D2) and (D3) are standard consequences of the primitive recursiveness of $\operatorname{Proof}_{\mathsf{RA}}$ (Lemma 3.4.3) and the representability of primitive recursive functions in $\mathsf{RA}$ (Proposition 1.3.3): $\mathsf{RA}$ can internally simulate concatenation of proof sequences and verify that the result is again a valid proof. See Boolos et al. [2012]. ∎
Theorem 3.5.2 (Diagonal Lemma).
@@
$\mathsf{RA}\;\vdash\;u\;\leftrightarrow\;\varphi(\ulcorner u\urcorner).$ (88)
By construction. Define the substitution function $\mathrm{sub}(e,n)$ to be the code of the sentence obtained by substituting the numeral $\overline{n}=S^{n}(0)$ for the free variable in the formula coded by $e$. The function $\mathrm{sub}$ is primitive recursive, it acts by bounded structural replacement in the formula tree, and is therefore representable in $\mathsf{RA}$ by Proposition 1.3.3. Let
$\psi(x):\equiv\varphi\big(\mathrm{sub}(x,x)\big)\quad\text{with}\quad e:=\ulcorner\psi\urcorner$ (89)
@@
$u\;\equiv\;\psi(\overline{e})\;\equiv\;\varphi\big(\mathrm{sub}(\overline{e},\overline{e})\big)\;\equiv\;\varphi(\ulcorner u\urcorner),$ (90)
@@
4 From Proof Checking to Constraints
@@ 4.1 The Hilbert–Ackermann Kernel
Remark 4.1.1.
We fix an automated proof system for $\mathsf{RA}$ in the style of Hilbert and Ackermann [1938]. Formulas are built from $\mathcal{L}_{\mathsf{RA}}$ atoms, implication ($\to$), and falsity ($\bot$). A proof is a finite sequence of formulas, each either a logical axiom instance ($\mathsf{K,S,EFQ}$), an instance of $\mathsf{RA}$ (Definition 1.1.2), or a consequence of modus ponens from earlier lines, concluding at the target. A total boolean checker decides correctness; being primitive recursive, it arithmetizes directly via the preceding encoding mechanisms. We then reflect each verification condition into polynomial equalities over $\mathbb{N}$, obtaining a constraint system whose solvability is equivalent to checker acceptance (Theorem 4.7.3) with every emitted equation of degree at most $3$ (Proposition 4.7.4).
Definition 4.1.2 (Formula syntax).
@@
$\mathsf{Form}\;::=\;\bot\mid(\varphi\to\psi),$ (91)
Atomic $\mathcal{L}_{\mathsf{RA}}$-formulas (equalities and their negations) and bounded quantifiers are encoded as nested $\bot/{\to}$ trees via the Gödel Numbering of Section 3.4; the two constructors suffice to represent the full $\mathcal{L}_{\mathsf{RA}}$-formula language. The mechanized kernel ($\mathtt{R04}$) uses exactly these two constructors.
Definition 4.1.3 (Axiom schemas).
The proof system admits logical and non-logical axiom schemas. The logical schemas K, S, and EFQ [Troelstra and Schwichtenberg, 2000] hold for all formulas $\varphi,\psi,\chi\in\mathsf{Form}$:
@@
    $\big(\varphi\to(\psi\to\chi)\big)\to\big((\varphi\to\psi)\to(\varphi\to\chi)\big).$ (93)
@@
    $\bot\to\varphi.$ (94)
  4. (iv)
    Non-logical axioms of $\mathsf{RA}$. Each axiom of $\mathsf{RA}$ (Definition 1.1.2)—the successor axioms (2)–(6), instances of the $\Delta_{0}$-induction schema, and instances of the $\mathrm{B}\Sigma_{1}$-collection schema—is accepted as a valid proof line when its $\mathsf{Form}$-encoding matches the corresponding pattern. Recognition of each schema instance is primitive recursive.
Definition 4.1.4 (Proof sequence).
@@
$p=(\varphi_{0},\;\ldots,\;\varphi_{n-1})$ (95)
@@
      Axiom: $\varphi_{i}$ is an instance of one of the axiom schemas (Definition 4.1.3).
@@
      $\varphi_{k}=(\varphi_{j}\to\varphi_{i}).$ (96)
@@
    $\varphi_{n-1}=\varphi.$ (97)
Definition 4.1.5 (Proof checker).
@@
$\mathrm{check}:\mathsf{ProofSeq}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\}$ (98)
that returns $\mathrm{true}$ if and only if $p$ satisfies the conditions of Definition 4.1.4 with target $\varphi$. The search for modus ponens witnesses at each line $i$ is bounded by $i$, so $\mathrm{check}$ is total and primitive recursive.
Theorem 4.1.6 (Kernel soundness).